OpenStudy (anonymous):

PLEASE HELP Write the equation in slope-intercept form of the line satisfying the following. m= -1/3, y-intercept (0,1) The equation in slope intercept form is

OpenStudy (anonymous):

IMPORTANT LINE RELATED EQUATIONS TO KNOW AND MEMORIZE slope formula m= slope/ gradiant -- same thing \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept -- in the form of (0,y) \[y=mx+b\]

OpenStudy (anonymous):

@jmallis Just plug values into the slope- intercept formula :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

"Ok" means you'll show your result HERE?

OpenStudy (anonymous):

It's y=-1/3x +1

OpenStudy (anonymous):

i see no problems with taht answer

OpenStudy (anonymous):

@jmallis Feel good yourself :)